# Collaborating with AI Agents: Field Experiments on Teamwork, Productivity, and Performance

### v1

@@ 4 Discussion
4.1 Theoretical, practical, and methodological contributions
Theoretical contributions.
This study advances theories of teamwork and collaboration by demonstrating how AI agents, particularly GPT-based models, reshape communication dynamics and workload distribution in human-AI teams (Schneider et al., 2021). We show that collaboration with GPT agents emphasizes task-oriented communication over social interactions, reducing social coordination costs and enabling participants to focus more on content generation. These findings extend theories of fit and complementarity in teamwork by illustrating how AI traits when aligned with human personality traits, enhance productivity and creativity (Ancona and Caldwell, 1992). For example, high-openness AI paired with open individuals fosters idea generation, while agreeable AI paired with extraverted individuals balances task-oriented and social communication. Tuning how AI agents interact and communicate is especially critical since communication patterns can predict team viability (Cao et al., 2021).
Additionally, this research contributes to our understanding of collaborative outcomes for GPT-based agents by highlighting their dual role as content generators and collaborators. Unlike prior work that frames AI tools as passive assistants (Noy and Zhang, 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024), our findings reveal that GPT-based agents actively influence the structure and quality of collaborative outputs, particularly in multimodal tasks. This study emphasizes the need to design GPT-based agents not only for efficiency but also for alignment with human workflows, paving the way for more nuanced and effective human-AI partnerships.
Practical contributions.
This study provides actionable insights for organizations and designers deploying AI in collaborative environments. First, we demonstrate that GPT-based agents can significantly reduce social coordination costs, allowing individuals to focus more on task completion and content generation. By streamlining communication and reducing the need for social or emotional exchanges, AI agents enable greater individual productivity, particularly for low-performing participants. This finding is especially relevant for teams with varying skill levels or in high-pressure tasks where minimizing coordination overhead is crucial.
Second, our results highlight the importance of tailoring AI agents to align with the personality traits of their human collaborators. For instance, pairing high-openness AI with open individuals and agreeable AI with extraverted partners improves productivity and performance. These insights offer a practical roadmap for organizations to design and deploy AI agents that enhance team performance by optimizing personality fit. While studies have shown heterogeneity in the positive productivity effects of generative AI across role, function, and organization (Jaffe et al., 2024), it was unclear whether and how different versions of AI agents influence productivity and for what types of people. Customizing AI behavior based on human traits can drive better collaboration outcomes, particularly in creative or iterative workflows.
Third, the study underscores the strengths and limitations of GPT-based agents in multimodal workflows. While these agents excel at enhancing text quality, they underperform in tasks involving image-related outputs. This suggests that organizations should pair GPT-based agents with complementary tools specifically designed for image generation and evaluation. Such integrations can mitigate the trade-offs observed in multimodal tasks, ensuring consistent quality across all aspects of output.
Finally, this work offers guidance for structuring collaborative workflows in mixed human-AI teams. By identifying how GPT agents influence communication patterns and workload distribution, we provide a framework for assigning tasks that maximize individual and team productivity. For example, GPT agents can handle bulk or iterative edits while human collaborators focus on creative ideation and final quality checks. This division of labor can enhance efficiency while leveraging the unique strengths of both humans and AI, offering practical implications for industries ranging from marketing to product development and creative design.
Methodological contributions.
This study introduces several methodological innovations that enhance our understanding of human-AI collaboration. First, the randomization of Human-Human and Human-AI teams represents a unique experimental design, allowing us to identify the distinct characteristics of human-human teamwork compared to collaborations involving AI agents. This approach provides causal evidence on how AI agents reshape communication patterns, workload distribution, and performance outcomes, filling a critical gap in prior research that often lacks such controls.
Second, we developed and employed a novel experimental ethnography approach, combining quantitative metrics with rich qualitative insights. The level of granularity in the data—capturing every time-stamped message, edit, and API call—enables a detailed reconstruction of collaboration workflows. This method bridges traditional experimental research with ethnographic approaches, offering a comprehensive view of team dynamics in real-time collaborative settings.
Third, this study introduces randomized personality prompts for AI agents, a methodological advance with significant implications for the burgeoning field of prompt engineering. Unlike traditional prompt engineering, which often focuses on one-shot outcomes, our approach explores how prompts influence interactions over time and within dynamic workflows (Schulhoff et al., 2024). By randomizing Big Five personality traits in AI agents, we provide a robust framework for studying the causal effects of personality alignment on collaboration outcomes. This methodology can be extended to test other dimensions of prompts, such as creativity, leadership styles, or domain expertise, across a range of interactive settings.
Finally, we built a state-of-the-art collaboration platform integrating the latest generative AI models and real-time collaborative tools, including APIs for ad editing and image generation. This platform not only advances the study of human-AI collaboration but also directly informs firms developing and deploying AI agents in workplace settings. By capturing how humans and AI interact in realistic, task-driven environments, the platform provides actionable insights for organizations aiming to design effective and efficient collaborative workflows.
@@ 4.2 Limitations and future directions
Several limitations in this study warrant further investigation. First, while we observe substantial differences in communication patterns between Human-Human and Human-AI teams, the underlying reasons remain unclear. These differences could stem from the nature of AI communication, which tends to prioritize content-related messages, or from human participants adjusting their behavior upon realizing their partner is an AI, reducing social or emotional exchanges. Future work is needed to elucidate these potential explanations and better understand the drivers of these shifts.
Second, the limitations of current AI models, particularly Vision GPT models, present challenges for multimodal tasks. These models are optimized for next-word generation and specific visual tasks, such as identifying items in images, but are not designed for nuanced assessments like image quality prediction. This limitation likely contributed to the lower image quality observed in Human-AI teams and underscores the need for purpose-built AI systems tailored to specific creative and evaluative tasks.
Finally, while our study provides a controlled experimental context, it may not fully capture the complexities of long-term collaboration with AI agents in real-world environments. Future research should explore how these dynamics evolve over extended periods and in more diverse task domains to validate and expand on our findings.
@@
This study opens several promising avenues for future research. First, while we randomized personality prompts to explore the interactive effects of human and AI personality alignment, future work could extend this approach to other dimensions of AI behavior. For example, prompts could be designed to vary in creativity, task orientation, or communication style, providing deeper insights into how different AI behaviors influence collaboration outcomes.
Second, the limitations of current Vision GPT models suggest a need for separate, specialized image generation models. Future studies could incorporate state-of-the-art visual models designed specifically for tasks like image quality prediction and generation, enabling a more nuanced understanding of AI contributions to multimodal creative tasks.
Third, our findings are grounded in the context of ad design, but they invite exploration across a range of other collaborative domains. Important contexts include software development (e.g., coding), data analysis, collaborative writing, and financial accounting. Investigating Human-AI collaboration in these areas could reveal domain-specific dynamics and inform the broader design of AI systems tailored to various professional workflows.
Finally, extending this research to longitudinal settings could provide insights into how Human-AI collaboration evolves over time. Long-term studies could examine the sustainability of productivity gains, the development of trust, and the potential for “learning effects” where humans adapt to working with AI agents or vice versa. These investigations would bridge the gap between controlled experimental settings and real-world applications, enhancing the generalizability of our findings.
@@ 4.3 Conclusion
This study provides novel insights into how AI agents reshape teamwork, productivity, and performance in collaborative settings. By introducing randomized personality prompts and leveraging a state-of-the-art collaboration platform, we demonstrate that Human-AI teams communicate more, focus on task-related content, and achieve higher individual productivity compared to Human-Human teams. While collaboration with AI agents enhances text quality, it introduces trade-offs in multimodal outputs like images. Importantly, the forthcoming field evaluation of ad click-through and view-through rates will shed light on how these findings translate to real-world outcomes. Together, our results highlight the transformative potential of AI agents in collaborative workflows while underscoring the importance of aligning their design with human traits and task requirements.
@@ C Prompts for AI ratings
D Full interaction effects for human and prompted AI personalities for the field study
Table D.1: Personality-heterogeneous prompt effects on CPC and CTR.
CPC CTR
Intercept 14.211*** 16.533*** -0.067 -0.141*
(2.920) (3.041) (0.057) (0.059)
Spend -0.144*** -0.150*** 0.007*** 0.008***
(0.019) (0.020) (0.000) (0.000)
Image -0.461** 0.005
(0.178) (0.003)
Text -0.314 0.012***
(0.169) (0.003)
Click 0.470* -0.006
(0.198) (0.004)
OpennessH -1.354 -1.425 0.048 0.051
(1.722) (1.718) (0.034) (0.033)
ConscientiousnessH -1.797 -1.913 0.088* 0.084
(2.300) (2.295) (0.045) (0.045)
ExtraversionH -3.003 -2.936 -0.023 -0.020
(1.849) (1.845) (0.035) (0.035)
AgreeablenessH 1.108 1.087 -0.038 -0.034
(2.082) (2.077) (0.041) (0.040)
NeuroticismH -1.295 -1.381 -0.002 -0.002
(1.906) (1.898) (0.037) (0.037)
OpennessAI 2.523 2.290 0.062 0.068
(2.268) (2.261) (0.044) (0.044)
ConscientiousnessAI -2.175 -1.972 -0.002 0.004
(2.315) (2.313) (0.046) (0.045)
ExtraversionAI -0.495 -0.583 -0.007 -0.004
(2.313) (2.304) (0.045) (0.045)
AgreeablenessAI -4.779* -4.457 0.027 0.013
(2.311) (2.312) (0.045) (0.045)
NeuroticismAI -0.911 -1.431 -0.017 -0.008
(2.327) (2.326) (0.046) (0.045)
OpennessH $\times$ OpennessAI -0.287 -0.082 -0.035 -0.044
(1.424) (1.428) (0.028) (0.028)
OpennessH $\times$ ConscientiousnessAI 0.525 0.494 0.025 0.027
(1.426) (1.420) (0.028) (0.028)
OpennessH $\times$ ExtraversionAI 2.151 2.095 -0.031 -0.033
(1.420) (1.416) (0.028) (0.027)
OpennessH $\times$ AgreeablenessAI 0.688 0.458 -0.010 -0.005
(1.388) (1.385) (0.027) (0.027)
OpennessH $\times$ NeuroticismAI -1.907 -1.645 -0.004 -0.008
(1.421) (1.420) (0.028) (0.028)
ConscientiousnessH $\times$ OpennessAI -0.695 -0.405 0.001 -0.002
(1.737) (1.734) (0.034) (0.034)
ConscientiousnessH $\times$ ConscientiousnessAI 1.083 0.931 -0.040 -0.041
(1.771) (1.766) (0.034) (0.034)
ConscientiousnessH $\times$ ExtraversionAI 1.315 1.477 -0.036 -0.035
(1.747) (1.743) (0.034) (0.034)
ConscientiousnessH $\times$ AgreeablenessAI 0.247 0.231 -0.029 -0.024
(1.728) (1.725) (0.033) (0.033)
ConscientiousnessH $\times$ NeuroticismAI 3.531* 3.537* -0.072* -0.068*
(1.764) (1.758) (0.034) (0.034)
ExtraversionH $\times$ OpennessAI -2.166 -2.394 0.017 0.025
(1.502) (1.498) (0.029) (0.029)
ExtraversionH $\times$ ConscientiousnessAI 1.178 1.116 -0.004 -0.009
(1.502) (1.503) (0.029) (0.029)
ExtraversionH $\times$ ExtraversionAI -0.836 -1.005 0.041 0.042
(1.522) (1.517) (0.029) (0.029)
ExtraversionH $\times$ AgreeablenessAI 2.991 3.105* -0.019 -0.020
(1.542) (1.537) (0.030) (0.029)
ExtraversionH $\times$ NeuroticismAI 0.588 0.483 0.055 0.056
(1.485) (1.480) (0.029) (0.028)
AgreeablenessH $\times$ OpennessAI -0.781 -0.788 -0.044 -0.046
(1.584) (1.578) (0.031) (0.031)
AgreeablenessH $\times$ ConscientiousnessAI 0.624 0.512 0.021 0.022
(1.572) (1.568) (0.031) (0.031)
AgreeablenessH $\times$ ExtraversionAI -2.511 -2.314 0.054 0.050
(1.619) (1.614) (0.032) (0.031)
AgreeablenessH $\times$ AgreeablenessAI 2.789 2.415 0.003 0.013
(1.625) (1.625) (0.032) (0.032)
AgreeablenessH $\times$ NeuroticismAI -1.877 -1.479 0.031 0.020
(1.590) (1.589) (0.031) (0.031)
NeuroticismH $\times$ OpennessAI -0.545 -0.655 -0.028 -0.025
(1.423) (1.418) (0.028) (0.027)
NeuroticismH $\times$ ConscientiousnessAI -0.579 -0.568 0.011 0.006
(1.450) (1.448) (0.028) (0.028)
NeuroticismH $\times$ ExtraversionAI 1.466 1.361 -0.016 -0.017
(1.486) (1.482) (0.029) (0.029)
NeuroticismH $\times$ AgreeablenessAI 1.530 1.570 0.016 0.021
(1.510) (1.510) (0.029) (0.029)
NeuroticismH $\times$ NeuroticismAI 0.956 1.156 0.059* 0.056*
(1.443) (1.439) (0.028) (0.028)
Campaign RE Yes Yes Yes Yes
Observations 1859 1859 2000 2000
Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Heteroskedasticity-robust standard errors are reported. AI{AI} refers to induced AI personality; $H$ refers to human personality covariates.
Table D.1: *

### v2

@@ 4 Discussion
4.1 Theoretical, managerial, and methodological implications
Theoretical implications.
This study advances theories of teamwork, collaboration, and human-AI interaction by demonstrating how GPT-based AI agents reshape work processes, communication dynamics, and task execution in collaborative settings (Schneider et al., 2021). Our findings reveal that AI agents shift communication from social and emotional exchanges to task-oriented interactions, reducing interpersonal coordination costs and enabling participants to prioritize content generation. This aligns with theoretical frameworks suggesting that AI can streamline team processes by automating routine or procedural tasks (Wilson and Daugherty, 2018), but it extends these models by uncovering trade-offs in creative, multimodal workflows. Specifically, while AI agents enhanced text quality in ad creation, they underperformed in image-related tasks, suggesting that theories of human-AI collaboration must account for task-specific competencies and limitations, particularly in domains requiring diverse cognitive and creative skills (Dell’Acqua et al., 2023).
Moreover, our work challenges the traditional conceptualization of AI as a passive tool or assistant (Noy and Zhang, 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024) by positioning GPT-based agents as active collaborators that shape the structure and quality of team outputs (Anthony et al., 2023, Makarius et al., 2020). For example, in our experiment, the AI’s proficiency in generating ad copy influenced participants to focus on generating text through instructions rather than direct edits. These changes in work processes suggest a need to reframe human-AI interaction theories to emphasize AI’s agency in collaborative processes. Rather than merely augmenting human efforts, AI agents co-create outcomes, introducing new dynamics that affect team performance.
Our findings also contribute to the literature on team cognition and distributed expertise (Collins et al., 2024). In human-human teams, expertise is often shared through social negotiation and emotional rapport, which fosters trust but can slow decision-making. In contrast, human-AI teams in our study bypassed much of this social overhead. This shift raises theoretical questions about how cognitive load and role allocation differ when humans collaborate with machines instead of other humans. Future theoretical work could explore these dynamics in other creative domains, such as product design, to develop a more nuanced understanding of how AI agents influence cognitive and collaborative processes across varied task environments.
Managerial implications.
Our work also offers actionable insights for managers and organizations seeking to integrate GPT-based AI agents into collaborative workflows, particularly in creative and productivity-driven environments. Our findings demonstrate that AI agents significantly reduce social coordination costs, allowing team members to focus on task execution and content generation. Managers in industries like marketing, content creation, or consulting can leverage AI agents to streamline routine tasks, such as drafting campaign materials or reports, freeing human employees for strategic activities like audience analysis or brand positioning.
However, the study also highlights critical limitations that managers must address to fully harness AI’s potential in multimodal workflows. While GPT-based agents excelled in text generation, they underperformed in image-related tasks. This trade-off supports the idea that organizations should adopt a hybrid approach along a jagged edge of AI capabilities (Dell’Acqua et al., 2023), allocating tasks within AI capabilities to agents and those outside AI capabilities, such as image selection, to humans. Such an approach can mitigate the quality disparities observed in our study and are particularly relevant in domains where compelling visual elements drive downstream outcomes.
Furthermore, our results provide a framework for structuring human-AI teams to optimize productivity and performance. AI’s ability to handle bulk editing tasks—as evidenced by the 71% fewer direct copy edits in human-AI teams—allowed participants to engage in more content ideation and quality control. Managers can apply this division of labor to design efficient workflows, assigning AI agents to repetitive or data-intensive tasks while reserving human expertise for creative oversight and final approvals. A practical example might involve a content production team where AI drafts blog posts or social media captions, which editors then polish for tone and brand alignment. This approach not only boosts efficiency but also leverages human strengths in contextual judgment, which AI currently lacks, particularly in nuanced visual tasks. To implement such workflows, organizations should invest in training programs that teach employees how to delegate tasks to AI effectively, fostering a collaborative culture that maximizes both technological and human contributions.
Finally, the study underscores the importance of aligning AI deployment with organizational goals and team dynamics. The comparable team-level productivity between human-AI and human-human teams, despite human-AI teams having half the human resources, suggests that AI can serve as a near-substitute for additional personnel in certain contexts. Managers could use this insight to scale team output without proportionally increasing headcount, a cost-effective strategy for startups or resource-constrained firms. However, they must remain mindful of the social dynamics altered by AI integration. The decrease in social and emotional messages in human-AI teams indicates less time spent on rapport-building, which could impact team cohesion in long-term projects. Managers should monitor team morale and consider hybrid models—such as combining AI assistance with periodic human-human collaboration—to maintain a balance between efficiency and interpersonal connection.
Methodological implications.
This study advances the methodological landscape for studying human-AI collaboration through the development of the Pairit platform, a novel experimental framework that redefines how researchers investigate team dynamics in real-time, task-driven settings. Unlike traditional approaches that rely on aggregate performance metrics or isolated AI interactions (Noy and Zhang, 2023, Dell’Acqua et al., 2023), Pairit captures granular, time-stamped data on every keystroke, message, edit, and API call, enabling a detailed reconstruction of collaborative workflows. In our experiment, this capability allowed us to analyze 183,691 messages, 1,960,095 text edits, and 63,656 image edits, revealing nuanced patterns such as how AI-driven suggestions shifted human effort from direct text editing to instructional messaging. This methodological advance allows us to analyze both aggregate and detailed dynamics of human-AI teamwork and offers researchers new insights into decision-making, task allocation, and interaction processes.
The randomization of team composition represents another significant methodological advance, moving beyond human-only versus human-AI comparisons to include human-human and human-AI pairings. Unlike prior research that often contrasts AI-augmented individuals with those working alone (Noy and Zhang, 2023, Vaccaro et al., 2024), our approach allowed us to evaluate the effects on collaboration itself, which is critical to understand given the proliferation of AI agents in workplaces. This randomization framework is also adaptable to other configurations, such as teams with multiple humans or multiple AI agents, enabling researchers to test how team composition influences outcomes in contexts like project management, educational group work, or distributed software development.
To fairly compare human-human and human-AI collaboration, another critical contribution is that AI agents are agentic. That is, AI agents on Pairit are capable of performing nearly all actions available to humans, including sending chat messages, editing text, selecting images, and generating new images via external APIs (e.g., DALL-E 3), with the sole exception of final ad submission in this study. This approach ensures direct comparability between human-AI and human-human conditions, a departure from prior studies where AI is often just a chatbot (Chen and Chan, 2024, Brynjolfsson et al., 2023). By enabling AI to act independently and actively, we analyze how AI’s active participation reshapes collaborative dynamics, which would not have been possible in experimental designs that treat AI as a passive tool.
Pairit’s multimodal capabilities further distinguish it as a methodological tool, supporting tasks that integrate text, images, and external API interactions. While our study focused on ad creation, Pairit’s architecture is extensible to diverse domains, such as software development (e.g., AI generating code while humans debug), scientific writing (e.g., AI drafting literature reviews), or multimedia content production (e.g., designing virtual reality environments). This versatility allows researchers to explore human-AI collaboration in tasks requiring varied cognitive and creative skills, addressing a gap in prior studies that often limit AI to text-based functions (Brynjolfsson et al., 2023). For example, a future study could use Pairit to analyze how AI agents assist in architectural planning, tracking both textual specifications and visual renderings to assess collaborative efficiency and output quality. By capturing multimodal interactions, Pairit provides a comprehensive framework for studying the interplay of different task components, a methodological leap forward for understanding complex workflows.
These methodological contributions have broader implications for experimental research and organizational practice. Pairit’s ability to simulate realistic collaborative environments offers a testing ground for refining theoretical models of human-AI interaction and informing the design of AI-driven workflows. For instance, researchers could leverage the platform to study how AI agents perform in agile software development sprints, using time-stamped data to optimize task assignments and evaluate productivity trade-offs. Ultimately, Pairit’s methodological innovations bridge the gap between controlled experiments and real-world applications and provides a robust, scalable framework for advancing our understanding of human-AI collaboration.
@@ 4.2 Limitations and future directions
While we conducted one of the first truly agent-randomized experiments in human-AI collaboration, combining laboratory and field tests, on a newly developed platform capable of recording fine-grained collaboration dynamics, our work is not without limitations. First, while we observe substantial differences in communication patterns between human-human and human-AI teams, the underlying mechanisms or explanations for these differences are not always obvious. They could stem from the nature of AI communication, which tends to prioritize content-related messages, or from human participants adjusting their behavior upon realizing their partner is an AI. Our exit surveys found that human participants were good at predicting whether their partner was an AI (2.7 points higher on a 7-point Likert scale, $p<.001$), suggesting AI communication and human adaptation are both at work in these mechanisms. But, future work is needed to elucidate these potential explanations and better understand the drivers of these shifts. Second, the limitations of current AI models, particularly Vision GPT models, present challenges for multimodal tasks. These models are optimized for next-word generation and specific visual tasks, such as identifying items in images, but are not designed for nuanced assessments like image quality prediction. This limitation likely contributed to the lower image quality observed in human-AI teams and underscores the need for AI systems tailored to specific creative and evaluative tasks. However, AI is a moving target and image selection and design competency likely will improve over time. As with any study of AI, results may drift as AI changes. Moreover, while our study provides a controlled experimental context, it may not fully capture the complexities of long-term collaboration with AI agents in real-world environments. Future research should explore how these dynamics evolve over extended periods and in more diverse task domains to validate and expand on our findings.
Finally, our study lacks a human alone condition, which prevents us from directly assessing how human-AI teams compare to individuals working independently. Without this baseline, we cannot precisely isolate AI’s marginal contribution to productivity and the coordination costs associated with collaboration. Nevertheless, our findings have practical implications for organizations looking to integrate AI agents into collaborative workflows, particularly in tasks requiring creativity and coordination. Our work lays a foundation for understanding human-AI collaboration in complex tasks, and future research can build on this by incorporating a human-alone condition to further disentangle marginal productivity effects.
@@
Our work opens several promising avenues for future research. First, the limitations of current Vision GPT models suggest a need for separate, specialized image generation models. Future studies could incorporate state-of-the-art visual models designed specifically for tasks like image quality prediction and generation, enabling a more nuanced understanding of AI contributions to multimodal creative tasks. Second, our findings are grounded in the context of ad design, but they invite exploration across a range of other collaborative domains. Important contexts include software development (e.g., coding), data analysis, collaborative writing, and financial accounting. Investigating Human-AI collaboration in these areas could reveal domain-specific dynamics and inform the broader design of AI systems tailored to various professional workflows. Finally, extending this research to longitudinal settings could provide insights into how Human-AI collaboration evolves over time. Long-term studies could examine the sustainability of productivity gains, the development of trust, and the potential for “learning effects” where humans adapt to working with AI agents or vice versa. These investigations would bridge the gap between controlled experimental settings and real-world applications, enhancing the generalizability of our findings. We intend to make Pairit freely available to academics advancing peer-reviewed research into human-AI collaboration. We believe this platform could support hundreds, if not thousands, of new experimental studies on the effects of AI on productivity, performance, and work processes.
@@ 4.3 Conclusion
This study provides novel insights into how AI agents reshape work processes, productivity, and performance in collaborative settings. Our findings reveal that collaborating with AI agents significantly increases communication while decreasing (but not completely eliminating) direct engagement with the work product. Human-AI teams focus more on content and process-oriented communication, reducing the need for social and emotional exchanges compared to human-human teams. Humans in human-AI teams experienced 73% greater productivity per worker, suggesting that AI agents can act as near-substitutes for additional human collaborators. However, AI had countervailing effects on different dimensions of output quality—while ad text quality improved in human-AI teams, image quality declined, highlighting a trade-off in multimodal tasks. Field evaluations of ad effectiveness showed that ads with higher image quality (produced by human-human collaborations) and higher text quality (produced by human-AI collaborations) performed significantly better on click-through rates, view through rates, and cost per click metrics.
A key innovation of this study is the introduction of Pairit, an experimental platform that enables real-time collaboration between humans and multimodal AI agents. By providing researchers with a controlled environment to study these dynamics, Pairit opens new avenues for understanding how AI agents can enhance teamwork and productivity. Its versatility allows for the exploration of a wide range of collaborative tasks beyond ad creation, including software development, data analysis, and collaborative writing, making it a valuable tool for researchers across disciplines. As AI agents become increasingly integrated into modern workflows, platforms like Pairit will be essential for advancing our knowledge of human-AI collaboration and designing more effective human-AI partnerships.